# WLM: Dynamics of an isolated Dwarf Irregular Galaxy Under Ram Pressure in the Local Group

### v1

@@
WLM is an archetypal dwarf irregular galaxy that has not experienced interactions with major Local Group galaxies within the past 8 Gyr. It has recently been shown that WLM is losing its gas due to ram pressure forces exerted by the surrounding intergalactic medium (IGM). In this work, we explore how ram pressure may also affect the WLM gas kinematics, and we show that its dynamics is especially perturbed at its outskirts, explaining the asymmetric rotation between the approaching and receding sides. Moreover, we have been able to decompose WLM in two main components, a compact one with a solid-body rotation that resembles a bar-like structure, and a more extended one with a characteristic double-horn profile suggesting an edge-on disk. The former is relatively unaffected by ram pressure while the latter has its dynamics considerably affected by ram pressure. This study shows that mass estimates of a dwarf galaxy like WLM should account for a full modeling of its dynamical components, especially accounting for its asymmetric rotation curve.
@@
The Wolf–Lundmark–Melotte (WLM) galaxy is an isolated gas rich dwarf irregular galaxy in the Local Group (LG) roughly 900 kpc away from the Milky Way and M31 (McConnachie, 2012). Within 1 Mpc, WLM is one of the dwarf galaxies that has not experienced interactions with the Milky Way or with M31 (Higgs et al. 2021). Owing to this spatial isolation throughout its evolutionary history, WLM has been a subject of extensive studies and has been characterized in the literature as a ”prototypical irregular dwarf”. In the hierarchical merger scenario of galaxy formation, dwarfs like WLM have merged to form larger galaxies like the Milky Way and M31. Hence studying an isolated dwarf like WLM gives us an insight into the properties of these progenitor dwarfs. Its high inclination has also made it a good candidate for dynamical studies (Hunter et al. 2012). However, the inclination poses a challenge to determine features within the thick disk, which has also been described as an oblate spheroid.
WLM therefore has a rich history of being observed by flagship telescopes across frequencies. Recent JWST+HST studies have probed its star formation history (McQuinn et al. 2024; Albers et al. 2019), revealing steady star formation for the past 7 billion years, owing to its gas rich nature. Molecular clouds contributing to this star formation traced by CO cores have also been extensively studied by ALMA (Archer et al. 2022) and in UV (Mondal et al. 2018). HI Radio observations have been performed with the VLA, GBT and the MeerKAT telescopes. They have provided us with an exquisite view into the neutral gas dynamics of WLM. For over a decade it has been noted in the literature, first by Kepley et al. (2007), that WLM shows a lopsided rotation curve on the approaching side of the galaxy. This triggered work to explain the asymmetry by considering a perturbed halo (Khademi et al. 2021), but this is not supported by the absence of past interactions for WLM.
Yang et al. (2022) found evidence of ram pressure stripping in the galaxy with the discovery of 4 trailing clouds detected by MeerKAT in the direction opposite of the galaxy’s proper motion which was obtained from Gaia DR3 by Battaglia et al. (2022). They were also able to show that the clouds had no stars associated with them, confirming that the gas has been stripped through ram pressure. Yang et al. (2022) were able to replicate the orientation and mass of these stripped clouds using hydrodynamical simulations. They considered multiple models of WLM through various densities of the IGM and assuming varying WLM velocities with respect to the IGM, to constrain its density in the surroundings of the LG. McConnachie et al. (2007) have reported a similar occurrence of gas loss due to ram pressure induced by the IGM within the LG for the Pegasus dwarf galaxy (DDO 216), which is also an isolated body.
@@
A large reservoir of baryonic matter in the universe is expected to be present in IGM filaments. Observational regimes ranging from Fast Radio Bursts to Lyman-alpha based methods have been used to put constraints on the cosmic IGM density (Connor et al. 2025; Jin et al. 2023). Intra-cluster medium of density - in large clusters has been studied and quantified using X-ray telescopes (Li et al. 2024). However, it is much sparser in less massive systems such as the LG, making it very hard to detect.
HI gas disks extend much farther beyond the stellar disks, especially for dwarf irregulars, and therefore, can be used to probe the total mass content of a galaxy to larger radii (Carignan and Freeman 1988; de Blok et al. 2008; Patra et al. 2016), under the assumption of dynamical equilibrium. In the case of WLM, an extreme dark matter (DM)-to-baryon ratio of 90:1 has been reported (Read et al. 2017). However, since the pioneering Gunn & Gott (1972) study, it is well established that the motion of galaxies through intragroup or intracluster medium perturbs their kinematics, particularly at the outer edges of their gaseous discs. Such interactions can drive the gas out of equilibrium, thereby biasing rotation curves and leading to systematic errors in the estimation of galactic masses. Therefore, having a thorough understanding of the DM halo extension and of the effect of the IGM on gas rich bodies is necessary to estimate the total mass content of dwarf irregulars.
There is growing evidence that ram pressure stripping due to the CGM of the Milky Way plays a major role in removing gas and transforming the dynamics of the Milky Way dwarfs (Hammer et al. 2024; Wang et al. 2024). Gas loss in WLM could also affect its dynamical evolution. The steady stripping of gas could also contribute to the two broad historical sub-classifications of isolated irregular dwarfs, those being gas rich and the other being mostly devoid of gas. Therefore, the goal of this paper is to use insights from state-of-art high velocity resolution 21cm radio observations of WLM from the MeerKAT telescope to study the kinematics for WLM and the effect of the IGM on gas dynamics.
In section 2 we revisit the Ianjamasimanana et al. (2020) MeerKAT 16 dish data to characterize the stripping gas more robustly and elaborate on the data reduction of the new MeerKAT 64 dish observations of WLM. In section 3, we look at the gas distribution in WLM and present a two-component Sersic fit. In section 4 we derive a rotation curve of WLM and discuss its dynamical properties. Section 5 elaborates on the kinematical decomposition in two components that we have performed on MeerKAT64 data, and in section 6 we discuss the consequences of our findings.
@@ 2.1 Revisiting the MeerKAT 16 dish Observations
Yang et al. (2022) used the observations from Ianjamasimanana et al. (2020), where they generated a mask by blanking pixels with values less than 3 times the RMS noise in the cube and used the MAFIA task in the MIRIAD software suite, leading them to detect the 4 stripped clouds. Motivated by trying to make a more robust detection of the stripped gas, we re-analysed the data cube with SoFiA-2 HI source finding tool. SoFiA-2 is optimised for HI surveys and can take into account noise variations and handles artifacts more reliably. We use the Smooth + Clip finder which iteratively smooths the cube with a defined set of spatial and spectral smoothing kernels, measuring the noise level at each smoothing scale and adding all pixels with an absolute flux density above a relative threshold (Serra et al. 2015; Westmeier et al. 2021a). We set the source finding threshold to be 4 times RMS, and use spatial smoothing kernels of sizes 3, 6, 9 , 12, 15, 18 pixels; as 3 pixels is just under the beam size and the smallest of the clouds detected by Yang et al. (2022) is roughly 18 pixels wide.
With the new SoFiA-2 mask applied to the data cube, we derive moment maps presented in Fig. 1. Instead of 4 isolated stripped clouds, we detect a continuous tail of stripped gas, akin to jellyfish galaxies (Roberts et al. 2021). For the main body of the galaxy, in the same region defined by Yang et al. (2022), we recover their total HI mass value of . However, Yang et al. (2022) have found a mass of for the stripped gas, while we recover an improved .
@@ 2.2 MeerKAT 64 Observation
Neutral hydrogen (H i) observations of the WLM galaxy were obtained with the MeerKAT radio telescope (Jonas 2018) as part of project SCI-20220822-BN-01, over two observing epochs on 2023 July 2–3 and July 3–4 (UTC). Observations were conducted in L-band 32k mode, which provides 32,768 frequency channels across an 856 MHz total bandwidth. For the purposes of H i spectral line imaging, a subset of 701 channels centered on the rest frequency of the H i line (1420.4058 MHz) was extracted, corresponding to a 2.289 MHz bandwidth. This yields a velocity coverage of km s-1 at a spectral resolution of 0.7 km s-1 per channel.
@@
To map the H i distribution in and around WLM, we used an 11-pointing mosaic covering a rectangular area of . The mosaic was arranged asymmetrically and extended toward the northwest, specifically designed to encompass several H i clouds previously detected in lower-resolution MeerKAT-16 observations. This layout provided complete coverage of both the central disk and the surrounding low-column-density H i environment. Each pointing was observed for a total integration time of approximately one hour. The primary flux and bandpass calibrator J1939–6342 was observed every 3 hours, while the complex gain (phase) calibrator J0025–2602 was observed approximately every 30 minutes.
@@
The data was processed on a virtual machine of the Spanish Prototype of the Square Kilometre Array (SKA) Regional Center (espSRC, Garrido et al. 2021). Both observing epochs were calibrated using the Containerised Automated Radio Astronomy Calibration (CARACal) pipeline (Józsa et al. 2020). CARACal provides an environment to carry out standard calibration and reduction steps, including RFI flagging, cross-calibration, self-calibration, continuum subtraction, and imaging.
@@
H i spectral line imaging was performed individually for each pointing using WSClean (Offringa et al. 2014) as part of the CARACal pipeline. We followed a two-step iterative strategy to produce the final image cubes. In the first step, a low-resolution cube was created using a Gaussian uv-taper of 90′′ and brings Briggs robust weighting of 1.5, to enhance sensitivity to extended low-surface-brightness emission. These cubes were cleaned using WSClean’s automatic masking at a 5 threshold, and the resulting masks were refined using SoFiA-2 (Westmeier et al. 2021b) to isolate real emission structures. In the second step, the refined masks were applied to produce final cubes at higher resolution using a uv-taper of 50′′. These final cubes were cleaned down to 0.5 per channel, ensuring both completeness of faint emission and suppression of sidelobe artefacts.
@@
The final data product is a linear mosaic composed of 11 pointings, each with a channel width of 0.7 km s-1 and a rms noise per channel of approximately 0.0034 Jy beam-1, measured before primary beam correction. The mosaic was produced using CARACal, with primary beam correction applied during the mosaicing process. The resulting synthesised beam size is 64′′ 60′′, and the average rms noise in line-free regions of the central, primary-beam-corrected mosaic area is approximately 0.0018 Jy beam-1 per 0.7 km s-1 channel. This imaging configuration was chosen to match the synthesized beam size of prior MeerKAT-16 observations, enabling direct comparison across datasets and resolutions.
@@
To account for beam variations among individual pointings across the mosaic, we estimated a relative scatter in beam area of 3.78%, which leads to a maximum flux density uncertainty of approximately 7.6%. Considering this uncertainty and the achieved sensitivity, the mosaic provides a 3 H i column density sensitivity of approximately cm-2 over a 20 km s-1 linewidth in the central region. The observation setup and relevant parameters are summarized in Table 2.
@@
The data was imaged with using natural weighting, several robust values, uv-tapers, and multi-scale deconvolution before settling on the above. The stripped gas detected in the earlier MEERKAT-16 observations was enabled by the compact 16-antenna configuration, which provided dense sampling at short baselines. In the 61-dish configuration used for our mosaic, the relative number of short spacings is significantly reduced, while the number of long baselines increases. This improves angular resolution but reduces sensitivity to large-scale, low-surface-brightness emission. We do not detect extended stripped faint structures with the 61 dish configuration, notably even features from the Magellanic Stream that appear clearly in single dish GASS observations (McClure-Griffiths et al. 2009) in the same velocity range of the mosaic field remain undetected, indicating that the limitation is intrinsic to the UV coverage rather than the imaging strategy. A full recovery of the diffuse tail would require combining our data with single-dish observations or acquiring targeted compact-configuration interferometric data. As this lies outside the scope of the present work, we plan to address the extended cloud in a dedicated follow-up study. While the reduced short-baseline sensitivity prevents recovery of these structures, including the diffuse tail, the 61-dish mosaic provides a high spectral resolution view of WLM’s main body and the spatial resolution necessary for a detailed kinematic analysis. The final mosaic is once again processed with SoFiA-2 to generate a mask with the smooth+clip finder spatial kernels set to 0,3,6 and 0, 3, 7, 15 for the spectral axis with the cubelets feature turned on, which gives us individual data cubes for each continuous source. A threshold of 5 RMS was set for detection. The cubelet that detected the main body of the galaxy was chosen as the data source, which was used in the study. The generation of moment maps, column densities, flux and HI mass reported in Fig. 2 and elsewhere in this study follow the same procedure and formulae reported in Ianjamasimanana et al. (2020).

### v3

@@
The Wolf–Lundmark–Melotte (WLM) galaxy is an archetypal dwarf irregular galaxy that has not experienced interactions with major Local Group galaxies within the past 8 Gyr. It has recently been shown that WLM is losing its gas due to ram pressure forces exerted by the surrounding intergalactic medium. In this work, we explored how ram pressure can also affect the WLM gas kinematics, and we show that its dynamics is especially perturbed at its outskirts, which explains the asymmetric rotation between the approaching and receding sides. Moreover, we have been able to decompose WLM into two main components, a compact one with a solid-body rotation that resembles a bar-like structure, and a more extended one with a characteristic double-horn profile suggesting an edge-on disc. The former is relatively unaffected by ram pressure while the dynamics of the latter is considerably affected. This study shows that mass estimates of a dwarf galaxy like WLM should involve a full modelling of its dynamical components, especially its asymmetric rotation curve.
@@
The Wolf–Lundmark–Melotte (WLM) galaxy is an isolated gas-rich dwarf irregular galaxy in the Local Group (LG) roughly 900 kpc away from the Milky Way and M31 (McConnachie, 2012). WLM is one of the few dwarf galaxies within 1 Mpc that have not experienced interactions with the Milky Way or with M31 (Higgs et al. 2021). Owing to this spatial isolation throughout its evolutionary history, WLM has been a subject of extensive studies and has been characterised in the literature as a ’prototypical irregular dwarf’. In the hierarchical merger scenario of galaxy formation, dwarfs like WLM have merged to form larger galaxies like the Milky Way and M31. Hence studying an isolated dwarf like WLM gives us an insight into the properties of these progenitor dwarfs. Its high inclination has also made it a good candidate for dynamical studies (Hunter et al. 2012). However, the inclination poses a challenge to determining features within the thick disc, which has been described as an oblate spheroid.
WLM therefore has a rich history of being observed by flagship telescopes across frequencies. Recent combined JWST and Hubble Space Telescope studies have probed its star formation history (McQuinn et al. 2024; Albers et al. 2019), revealing steady star formation for the past 7 billion years owing to its gas-rich nature. Molecular clouds contributing to this star formation traced by CO cores have also been extensively studied by the Atacama Large Millimeter/submillimeter Array (ALMA) (Archer et al. 2022) and in UV by the Ultraviolet Imaging Telescope aboard ASTROSAT (Mondal et al. 2018). HI Radio observations have been performed with the Very Large Array (VLA), Green Bank Telescope (GBT) and MeerKAT telescopes. They have provided us with an exquisite view into the neutral gas dynamics of WLM. For over a decade it has been noted in the literature, first by Kepley et al. (2007), that WLM shows a lopsided rotation curve on the approaching side of the galaxy. This triggered work to explain the asymmetry by considering a perturbed halo (Khademi et al. 2021), but the absence of past interactions disfavours this hypothesis.
Yang et al. (2022) found evidence of ram pressure stripping in the galaxy with the discovery of four trailing clouds detected by MeerKAT in the direction opposite to the galaxy’s proper motion which was obtained from Gaia Data Release 3 by Battaglia et al. (2022). They were also able to show that the clouds had no stars associated with them, confirming that the gas has been stripped through ram pressure. Yang et al. (2022) were able to replicate the orientation and mass of these stripped clouds using hydrodynamical simulations. They considered multiple models of WLM through various densities of the intergalactic medium (IGM) and assuming varying WLM velocities with respect to the IGM, to constrain its density in the surroundings of the LG. McConnachie et al. (2007) reported a similar occurrence of gas loss due to ram pressure induced by the IGM within the LG for the Pegasus dwarf galaxy (DDO 216), which is also an isolated body.
@@
A large reservoir of baryonic matter in the Universe is expected to be present in IGM filaments. Observational regimes ranging from fast-radio-bursts to Lyman-alpha-based methods have been used to place constraints on the cosmic IGM density (Connor et al. 2025; Jin et al. 2023). An intra-cluster medium of density - in large clusters has been studied and quantified using X-ray telescopes (Li et al. 2024). However, it is much sparser in less massive systems such as the LG, making it very hard to detect.
HI gas discs extend much farther beyond the stellar discs, especially for dwarf irregulars, and therefore, can be used to probe the total mass content of a galaxy to larger radii (Carignan and Freeman 1988; de Blok et al. 2008; Patra et al. 2016), under the assumption of dynamical equilibrium. In the case of WLM, an extreme dark matter (DM)-to-baryon ratio of 90:1 has been reported (Read et al. 2017). However, since the pioneering Gunn & Gott (1972) study, it is well established that the motion of galaxies through intra-group or intra-cluster medium perturbs their kinematics, particularly at the outer edges of their gaseous discs. Such interactions can drive the gas out of equilibrium, thereby biasing rotation curves and leading to systematic errors in the estimation of galactic masses. Therefore, having a thorough understanding of the DM halo extension and the effect of the IGM on gas-rich bodies is necessary to estimate the total mass content of dwarf irregulars.
There is growing evidence that ram pressure stripping due to the circumgalactic medium of the Milky Way plays a major role in removing gas and transforming the dynamics of the Milky Way dwarfs (Hammer et al. 2024; Wang et al. 2024). Gas loss in WLM could also affect its dynamical evolution. The steady stripping of gas could also contribute to the two broad historical sub-classifications of isolated irregular dwarfs, those being gas-rich and the other being mostly devoid of gas. Therefore, the goal of this paper is to use insights from state-of-art high-velocity-resolution 21cm radio observations of WLM from the MeerKAT telescope to study the kinematics of WLM and the effect of the IGM on gas dynamics.
In section 2 we revisit the Ianjamasimanana et al. (2020) MeerKAT 16-dish (MeerKAT-16 hereafter) data to characterize the stripping gas more robustly and elaborate on the data reduction of the new MeerKAT 64-dish (MeerKAT-64) observations of WLM. In section 3, we look at the gas distribution in WLM and present a two-component Sérsic fit. In section 4 we derive the rotation curve for WLM and discuss its dynamical properties. Section 5 elaborates on the kinematical decomposition that we performed on the MeerKAT-64 data, and in section 6 we discuss the consequences of our findings.
@@ 2.1 Revisiting the MeerKAT 16 dish observations
Yang et al. (2022) used the observations from Ianjamasimanana et al. (2020), where they generated a mask by blanking pixels with values less than 3 times the RMS noise in the cube and used the MAFIA task in the MIRIAD software suite, leading them to detect the four stripped clouds. Motivated by trying to make a more robust detection of the stripped gas, we re-analysed the data cube with the SoFiA-2 HI source finding tool. SoFiA-2 is optimised for HI surveys and can take into account noise variations and it handles artefacts more reliably. We used the Smooth + Clip finder which iteratively smooths the cube with a defined set of spatial and spectral smoothing kernels, measuring the noise level at each smoothing scale and adding all pixels with an absolute flux density above a relative threshold (Serra et al. 2015; Westmeier et al. 2021). We set the source finding threshold to be 4 times the RMS and used spatial smoothing kernels of sizes 3, 6, 9 , 12, 15, and 18 pixels; as 3 pixels is just under the beam size and the smallest of the clouds detected by Yang et al. (2022) is roughly 18 pixels wide.
With the new SoFiA-2 mask applied to the data cube, we derived the moment maps presented in Fig. 1. Instead of four isolated stripped clouds, we detect a continuous tail of stripped gas, akin to jellyfish galaxies (Roberts et al. 2021). For the main body of the galaxy, in the same region defined by Yang et al. (2022), we recover their total HI mass value of . However, Yang et al. (2022) have found a mass of for the stripped gas, while we recover an improved .
@@ 2.2 MeerKAT 64 Observation
Neutral hydrogen (H i) observations of the WLM galaxy were obtained with the MeerKAT radio telescope (Jonas 2018) as part of project SCI-20220822-BN-01, over two observing epochs on 2023 July 2–3 and July 3–4 (UTC). Observations were conducted in L-band 32k mode, which provides 32,768 frequency channels across an 856 MHz total bandwidth. For the purposes of H i spectral line imaging, a subset of 701 channels centred on the rest frequency of the H i line (1420.4058 MHz) was extracted, corresponding to a 2.289 MHz bandwidth. This yields a velocity coverage of km s-1 at a spectral resolution of 0.7 km s-1 per channel.
@@
To map the H i distribution in and around WLM, we used an 11-pointing mosaic covering a rectangular area of . The mosaic was arranged asymmetrically and extended towards the north-west, specifically designed to encompass several H i clouds previously detected in lower-resolution MeerKAT-16 observations. This layout provided complete coverage of both the central disc and the surrounding low-column-density H i environment. Each pointing was observed for a total integration time of approximately one hour. The primary flux and bandpass calibrator J1939–6342 was observed every 3 hours, while the complex gain (phase) calibrator J0025–2602 was observed approximately every 30 minutes.
@@
The data were processed on a virtual machine of the Spanish Prototype of the Square Kilometre Array (SKA) Regional Center(espSRC)(Garrido et al. 2021). Both observing epochs were calibrated using the Containerised Automated Radio Astronomy Calibration (CARACal) pipeline (Józsa et al. 2020). CARACal provides an environment to carry out standard calibration and reduction steps, including radio frequency interference flagging, cross-calibration, self-calibration, continuum subtraction, and imaging.
@@
H i spectral line imaging was performed individually for each pointing using WSClean (Offringa et al. 2014) as part of the CARACal pipeline. We followed a two-step iterative strategy to produce the final image cubes. In the first step, a low-resolution cube was created using a Gaussian uv-taper of 90′′ and brings Briggs robust weighting of 1.5, to enhance sensitivity to extended low-surface-brightness emission. These cubes were cleaned using WSClean’s automatic masking at a 5 threshold, and the resulting masks were refined using SoFiA-2 (Westmeier et al. 2021) to isolate real emission structures. In the second step, the refined masks were applied to produce final cubes at higher resolution using a uv-taper of 50′′. These final cubes were cleaned down to 0.5 per channel, ensuring both completeness of faint emission and suppression of sidelobe artefacts.
@@
The final data product is a linear mosaic composed of 11 pointings, each with a channel width of 0.7 km s-1 and a rms noise per channel of approximately 0.0034 Jy beam-1, measured before primary beam correction. The mosaic was produced using CARACal, with primary beam correction applied during the mosaicing process. The resulting synthesised beam size is 64′′ 60′′, and the average rms noise in line-free regions of the central, primary-beam-corrected mosaic area is approximately 0.0018 Jy beam-1 per 0.7 km s-1 channel. This imaging configuration was chosen to match the synthesised beam size of prior MeerKAT-16 observations, enabling direct comparisons across datasets and resolutions.
@@
To account for beam variations among individual pointings across the mosaic, we estimated a relative scatter in beam area of 3.78%, which leads to a maximum flux density uncertainty of approximately 7.6%. Considering this uncertainty and the achieved sensitivity, the mosaic provides a 3 H i column density sensitivity of approximately cm-2 over a 20 km s-1 linewidth in the central region. The observation setup and relevant parameters are summarised in Table 2.
@@
The data were imaged with using natural weighting, several robust values, uv-tapers, and multi-scale deconvolution before settling on the above. The stripped gas detected in the earlier MeerKAT-16 observations was enabled by the compact 16-antenna configuration, which provided dense sampling at short baselines. In the 61-dish configuration used for our mosaic, the relative number of short spacings is significantly reduced, while the number of long baselines increases. This improves angular resolution but reduces sensitivity to large-scale, low-surface-brightness emission. We do not detect extended stripped faint structures with the 61-dish configuration, notably even features from the Magellanic Stream that appear clearly in single dish Parkes Galactic All Sky Survey (GASS) observations (McClure-Griffiths et al. 2009) in the same velocity range of the mosaic field remain undetected, indicating that the limitation is intrinsic to the UV coverage rather than the imaging strategy. A full recovery of the diffuse tail would require combining our data with single-dish observations or acquiring targeted compact-configuration interferometric data. As this lies outside the scope of the present work, we plan to address the extended cloud in a dedicated follow-up study. While the reduced short-baseline sensitivity prevents recovery of these structures, including the diffuse tail, the 61-dish mosaic provides a high spectral resolution view of WLM’s main body and the spatial resolution necessary for a detailed kinematic analysis. The final mosaic is once again processed with SoFiA-2 to generate a mask with the smooth+clip finder spatial kernels set to 0,3,6 and 0, 3, 7, 15 for the spectral axis with the cubelets feature turned on, which gives us individual data cubes for each continuous source. A threshold of 5 RMS was set for detection. The cubelet that detected the main body of the galaxy was chosen as the data source, which was used in the study. The generation of moment maps, column densities, flux and HI mass reported in Fig. 2 and elsewhere in this study follow the same procedure and formulae reported in Ianjamasimanana et al. (2020).